Suppose we observe the following rates: 1R1 = 4.5%, 1R2 = 6.8%. If the unbiased expectations theory of the term structure of interest rates holds, what is the one-year interest rate expected one year from now, E(2r1)? (Do not round intermediate calculations. Round your answer to 2 decimal places.
Answer #1

1 + 1R2 = {(1 + 1R1)(1 + E(2r1))}1/2

1.068 = {1.045(1 + E(2r1))}1/2

1.140624 = 1.045(1 + E(2r1))

1.140624 / 1.045 = 1 + E(2r1)

1 + E(2r1) = 1.0915062201

E(2r1) = 0.0915 or 9.15%